# Streaming Max-Cut in General Metrics

### v1

@@
Streaming Max-Cut in General Metrics
Shaofeng H.-C. Jiang Center on Frontiers of Computing Studies, School of Computer Science, Peking University, Beijing, China. Email: shaofeng.jiang@pku.edu.cn. Partially supported by Beijing Natural Science Foundation (QY-24036).    Pan Peng School of Computer Science and Technology, University of Science and Technology of China (USTC), Hefei, China. Email: ppeng@ustc.edu.cn.    Haoze Wang School of Electronics Engineering and Computer Science, Peking University, Beijing, China. Email: 2200012915@stu.pku.edu.cn. Partially supported by Beijing Natural Science Foundation (QY-24036).
(October 6, 2025)
$\operatorname{Max-Cut}$ is a fundamental combinatorial optimization problem that has been studied in various computational settings. In this work, we initiate the study of its streaming complexity in general metric spaces with access to distance oracles. We give a $(1+\epsilon)$-approximation algorithm for estimating the $\operatorname{Max-Cut}$ value sliding-window streams using only poly-logarithmic space. This is the first sliding-window algorithm for $\operatorname{Max-Cut}$ even in Euclidean spaces, and it achieves a similar error-space tradeoff as the state-of-the-art insertion-only algorithms in Euclidean settings [Chen, Jiang, Krauthgamer, STOC’23], but without relying on Euclidean structures. In sharp contrast, we prove a polynomial-space lower bound for any $\operatorname{poly}(n)$-approximation in the dynamic streaming setting. This yields a separation from the Euclidean case, where the polylogarithmic-space $(1+\epsilon)$-approximation extends to dynamic streams.
@@ 1 Introduction
$\operatorname{Max-Cut}$ is a fundamental combinatorial optimization problem and it has been a central topic in algorithm study under various settings, including polynomial-time approximation algorithms [GW95, FDLV96, MS08, Yar14], streaming algorithms [FS05, CJK23, MW25], and sublinear-time algorithms [GGR98, ADLVKK03, RV07, Ind01, PY23]. In this paper, we initiate the study of the streaming complexity of metric $\operatorname{Max-Cut}$. In metric $\operatorname{Max-Cut}$, we are given an (underlying) general metric space $(V,\operatorname{dist})$ and a set of input data points $P\subseteq V$. The goal is to find a subset $S\subseteq P$ such that its cut value, defined as
@@
This streaming metric $\operatorname{Max-Cut}$ problem has been recently studied in the special case when the metric is Euclidean, and $(1+\epsilon)$-approximation has been obtained using small space [FS05, CJK23, MW25], even when the stream has both insertion and deletions. However, these results crucially exploit the Euclidean structure. A fundamental question is whether such structural assumptions are necessary for obtaining $(1+\epsilon)$-approximations in the streaming model, and whether the results can be extended to arbitrary metric spaces.
@@ 1.1 Our Results
In this work, we systematically address the streaming complexity of metric $\operatorname{Max-Cut}$, and this is the first set of results for streaming $\operatorname{Max-Cut}$ in general metrics. We focus on three standard streaming models: (1) Insertion-only streams: the stream consists solely of point insertions, and the dataset grows monotonically. (2) Sliding-window streams: the stream consists of point insertions, but only the most recent $w$ points ($w\geq 1$) are considered active, and the algorithm must maintain an estimate after every update. When $w=\infty$, this reduces to the insertion-only case. (3) Dynamic streams: the stream allows both insertions and deletions of points.
Our key conceptual message is that the streaming complexity demonstrates a strong separation between the insertion-only streams (and more generally sliding-window streams) and dynamic streams, by giving competitive streaming algorithms for sliding-window streams and a strong lower bound for dynamic streams. This also separates the Euclidean case from the general metric case, since in Euclidean spaces, $(1+\epsilon)$-approximation with poly-logarithmic space is known even for dynamic streams [FS05, CJK23, MW25].
@@
The above result is based on a new algorithm for metric $\operatorname{Max-Cut}$ in the insertion-only streams with polylogarithmic space (see Theorem 3.1). As mentioned, prior streaming bounds for metric $\operatorname{Max-Cut}$ were known only in Euclidean spaces, and the state-of-the-art [CJK23, MW25] was a $(1+\epsilon)$-approximation using space $\operatorname{poly}(\epsilon^{-1}d\log\Lambda)$ for points coming from $\Lambda]^{d}$ ($\Lambda\in\mathbb{N}$ and $\Lambda]:=\{1,\ldots,\Lambda\}$). Furthermore, it is unknown if this Euclidean bound holds for sliding window [^1] While the result of [CJK23] works for dynamic streams, their setting is not comparable to our sliding-window streams as the deletions must be explicitly given in their dynamic streaming setting whereas the sliding-window deletes an element implicitly. . Our result nearly matches this competitive trade-off, works for sliding window, and works for general metrics which does not rely on any specific structure of Euclidean spaces [^2] Our result readily applies to Euclidean case with an additional $O(d)$ factor in space complexity, since the trivial distance oracle that returns the $\ell_{2}$ distance between two points in $\mathbb{R}^{d}$ only uses $O(d)$ space. .
@@
Our second result is a lower bound for dynamic streams, and it shows that no algorithm using small space can achieve even $\operatorname{poly}(n)$-approximation in general metric spaces. Here, the input of a dynamic stream consists of both insertions and deletions of point (IDs), and the algorithm needs to report the (estimate of the) $\operatorname{Max-Cut}$ value at the end of the stream, with respect to the inserted points that are not yet deleted.
@@
There is no dynamic streaming algorithm that estimates the metric $\operatorname{Max-Cut}$ value within $\operatorname{poly}(n)$ factor using space $o(n^{1/3})$ with constant probability, where the $n$ is the size of the stream.
@@
In fact, our lower bound works under a stronger assumption about the distance oracle (which means this lower bound is actually stronger than the stated version which uses the default model as in Theorem 1.2), such that the algorithm is allowed to query the distance between any pair of points that ever appear in the stream, not only the ones that are currently stored by the algorithm which is the default model for our upper bound. We remark that apart from the default model and the model for our lower bound, a recent work [KPSW25] introduces an even stronger oracle model which allows the algorithm to query any pairs in the metric (not only those appear in the stream). Under this model, a lower bound for diameter problem is established [KPSW25], and it is interesting to establish the $\operatorname{Max-Cut}$ lower bound under the same model (which will be stronger than Theorem 1.2).
@@ 1.2.1 Upper Bounds
For insertion-only streams, our high-level idea is to (approximately) implement an importance sampling framework over general metrics, suggested in [CJK23] to devise a Euclidean streaming algorithm for $\operatorname{Max-Cut}$. In this framework, each data point $x\in P$ is sampled with probability roughly proportional to its total distance from all other points, denoted $q(x):=\sum_{y\in P}\mathrm{dist}(x,y)$. With a sample $S\subseteq P$ of size $\operatorname{poly}(\epsilon^{-1})$, we then evaluate $\operatorname{Max-Cut}(S)$. As in [CJK23] (restated in Lemma 3.8), this evaluation yields a $(1+\epsilon)$-approximation to $\operatorname{Max-Cut}(P)$. In our distance oracle model, evaluating $\operatorname{Max-Cut}(S)$ is straightforward since we can query the oracle to obtain all pairwise distances within $S$.
The main challenge lies in performing the importance sampling. To implement this, we propose the following reservoir-sampling style procedure. We note that reservoir sampling is a classic procedure for generating uniform samples, but here our novelty is that we adapt it to generate importance samples in a metric space whose importance score $q(x)$ is not given in advance. The process maintains a single sample $s\in P$. Label the data points in $P:=\{p_{1},\ldots,p_{n}\}$ according to their order of arrival in the stream, and let $P_{i}:=\{p_{1},\ldots,p_{i}\}$ denote the prefix of the first $i$ points. When the $i$-th point $p_{i}$ is inserted, instead of computing $q(p_{i})$, which depends globally on the entire dataset $P$, we consider the prefix-based quantity $q^{\prime}(p_{i}):=\sum_{x\in P_{i}}\operatorname{dist}(p_{i},x)$. Then we replace the current sample with $p_{i}$ with probability $\frac{q^{\prime}(p_{i})}{\sum_{j\leq i}q^{\prime}(p_{j})}$, and do nothing otherwise. This update rule is analogous to sampling proportional to $q(x)$, except that the probabilities are computed only with respect to the observed prefix.
In our concrete implementation, we can only maintain approximate values of $q^{\prime}(x)$, which introduces additional complications for both the algorithm and the analysis. Observe that $q^{\prime}(x)$ coincides with the $1$-median objective $\operatorname{med}(P_{i},x)$: for $Q\subseteq V$ and $y\in V$, we define $\operatorname{med}(Q,y):=\sum_{y^{\prime}\in Q}\operatorname{dist}(y^{\prime},y)$. It is known that an $\epsilon$-coreset (see Definition 3.3) for the $1$-median problem can be maintained efficiently in insertion-only streams (see Lemma 3.4). This result serves as a key building block for our final algorithm, and the $\epsilon$-approximation error is carefully accounted for in the analysis of Lemma 3.8.
@@
To leverage the framework of [BO10], we need to show that our insertion-only algorithm provides an approximation of a smooth function. Roughly speaking, a function $f$ is smooth, if for any $0<\epsilon<1$, there exist $0<\beta<\alpha<1$, such that if some $B\subseteq A\subseteq V$ satisfies $(1-\beta)f(A)\leq f(B)$ then $(1-\alpha)f(A\cup C)\leq f(B\cup C)$ for every $C\subseteq V$. A natural attempt is to prove the smoothness for $f=\operatorname{Max-Cut}$. However, somewhat surprisingly, $f(\cdot)=\operatorname{Max-Cut}(\cdot)$ does not satisfy this smoothness definition. To see this, consider the 1D case, and $f$ does not satisfy the condition for $A:=\{1,1\},B:=\{0\},C:=\{0\}$, because $f(A)=f(B)=0$ but $f(A\cup C)>0$ and $f(B\cup C)=0$.
We identify that the issue is caused by 0-valued $f$ (such as $f(B\cup C)=0$ as in the above example). Hence, our new idea is to add a tiny term to $\operatorname{Max-Cut}$, and define $f(S):=\operatorname{Max-Cut}(S)+\frac{\epsilon}{n}|S|$ to bypass this issue. We show that this new $f$ is indeed smooth. On the other hand, this tiny term does not add much error to $\operatorname{Max-Cut}$, so estimating $f$ suffices for estimating $\operatorname{Max-Cut}$. Indeed, if $\operatorname{Max-Cut}(P)>0$ (where $P$ is the dataset), then the added term is at most $\epsilon\operatorname{Max-Cut}(P)$, due to an assumption that the minimum non-zero distance is $1$. Hence, in this case, we can readily use the smooth histogram framework and reduce to estimating $f$ in an insertion-only stream. Finally, the other case of $\operatorname{Max-Cut}(P)=0$ can be easily detected in a sliding window and hence can be handled separately.
1.2.2 The Lower Bound
@@
At a high level, the metric of the hard input consists of many clusters such that the distance between the clusters is $\Delta\geq\Omega(n^{1/3})$ (where $n$ is the number of points), and inside the clusters the pairwise distance is $1$ (very small compared with $\Delta$), except for one special cluster of size $\Theta(n^{1/3})$, denoting whose point set as $S$. This special cluster $S$ has an outlier point whose distance to every other |S|1|S|-1 points in $S$ is some parameter $k\in\{1,\Delta\}$, thereby defining the two hard instances.
The stream first sequentially inserts all points, then delete every point except for the special cluster $S$, so this special cluster $S$ is the final input dataset. If $k=1$ then $\operatorname{Max-Cut}(S)\leq O(|S|^{2})<\Delta|S|$, whereas if $k=\Delta$ then $\operatorname{Max-Cut}(S)\geq\Delta|S|$. Furthermore, any algorithm with approximation ratio $o({\Delta}/{|S|})$ must be able to tell if $k=1$ or $k=\Delta$, and this is the main claim of our proof.
Without diving into the proof details, we discuss the intuition of why our input is hard in a conceptual sense. The first angle is how the deletions make it hard. Indeed, the $\operatorname{Max-Cut}$ for the entire input set (containing all points) is easy to approximate as ignoring the special cluster, or even ignoring any constant fraction of clusters, does not change the objective by much. However, the deletion forces the algorithm to learn the structure of a tiny special cluster $S$ accurately, which is a hard task. Another angle is how we utilize the general metric and distance oracle setting, especially how this is different from the Euclidean case. In fact, the metric in our hard input can be embedded into Euclidean spaces without large distortion, hence it is not the metric itself that is hard. What makes it harder than Euclidean case, is that the algorithm can only learn the distances through a distance oracle, without the access to the vector representation in Euclidean spaces. In particular, this forces the algorithm to explicitly query the distance between the outlier point and some other point in $S$ to differentiate whether $k=1$ or $k=\Delta$ (which is crucial for approximating the $\operatorname{Max-Cut}$), and explicitly discover the outlier point is difficult. On the other hand, in Euclidean spaces one can use structures such as (space efficient) tree embedding to approximately learn the cluster structure and differentiate the outlier, without explicitly knowing the outlier point (see e.g., [CJK23]).

### v2

@@
Abstract
$\operatorname{Max-Cut}$ is a fundamental combinatorial optimization problem that has been studied in various computational settings. We initiate the study of its streaming complexity in general metric spaces with access to distance oracles. We give a $(1+\epsilon)$-approximate algorithm for estimating the $\operatorname{Max-Cut}$ value in sliding-window streams using only poly-logarithmic space. This is the first sliding-window algorithm for $\operatorname{Max-Cut}$ even in Euclidean spaces, and it matches a known insertion-only space bound in the special case of Euclidean spaces [Chen, Jiang, Krauthgamer, STOC’23]. In sharp contrast, we give a $\operatorname{poly}(n)$-space lower bound in the dynamic streaming setting. This yields a separation from the Euclidean case, where the polylogarithmic-space $(1+\epsilon)$-approximation extends to dynamic streams.
@@ 1 Introduction
$\operatorname{Max-Cut}$ is a fundamental combinatorial optimization problem and it has been a central topic in algorithm study under various settings, including polynomial-time approximation algorithms [GW95, FDLV96, MS08, Yar14], streaming algorithms [FS05, CJK23, MW26], and sublinear-time algorithms [GGR98, ADLVKK03, RV07, Ind01, PY23]. In this paper, we initiate the study of the streaming complexity of metric $\operatorname{Max-Cut}$. In metric $\operatorname{Max-Cut}$, we are given an (underlying) general metric space $(V,\operatorname{dist})$ and a set of input data points $P\subseteq V$. The goal is to find a subset $S\subseteq P$ such that its cut value, defined as
@@
This streaming metric $\operatorname{Max-Cut}$ problem has been recently studied in the special case when the metric is Euclidean, and $(1+\epsilon)$-approximation has been obtained using small space [FS05, CJK23, MW26], even when the stream has both insertion and deletions. However, these results crucially exploit the Euclidean structure. A fundamental question is whether such structural assumptions are necessary for obtaining $(1+\epsilon)$-approximations in the streaming model, and whether the results can be extended to arbitrary metric spaces.
@@ 1.1 Our Results
In this work, we systematically address the streaming complexity of metric $\operatorname{Max-Cut}$, and this is the first set of results for streaming $\operatorname{Max-Cut}$ in general metrics. We focus on three standard models: (1) Insertion-only streams: the stream consists solely of point insertions. (2) Sliding-window streams: the stream consists of point insertions, but only the most recent $w$ points ($w\geq 1$) are considered active, and the algorithm must maintain an estimate after every update. When $w=\infty$, this reduces to the insertion-only case. (3) Dynamic streams: the stream allows both insertions and deletions of points. For both the insertion-only and dynamic streaming setting, the algorithm needs only output at the end of the stream.
Our key conceptual message is that the streaming complexity of $\operatorname{Max-Cut}$ demonstrates a strong separation between the insertion-only streams (and more generally sliding-window streams) and dynamic streams, by giving competitive streaming algorithms for sliding-window streams and a strong lower bound for dynamic streams. This also separates the Euclidean case from the general metric case, since in Euclidean spaces, $(1+\epsilon)$-approximation with poly-logarithmic space is known even for dynamic streams [FS05, CJK23, MW26].
As in the standard litereature of streaming algorithms, we measure the space in words, and each word can store $\operatorname{poly}\log N$ bits, where $N$ is the input size. We condiser the following distance oracle model throughout.
Definition 1.1 (Distance Oracle Model).
Each metric point is represented by a unique identifier (ID). The algorithm accesses distances only through a distance oracle, which takes two point IDs and returns their metric distance. The oracle representation itself is not charged to the algorithm’s space complexity, and we assume each ID or distance fits in one word.
@@
The above result is based on a new algorithm for metric $\operatorname{Max-Cut}$ in the insertion-only streams with polylogarithmic space (see Theorem 3.1). As mentioned, prior streaming bounds for metric $\operatorname{Max-Cut}$ were known only in Euclidean spaces, and the state-of-the-art [CJK23, MW26] was a $(1+\epsilon)$-approximation using space $\operatorname{poly}(\epsilon^{-1}d\log\Delta)$ for points coming from $\Delta]^{d}$ ($\Delta\in\mathbb{N}$ and $\Delta]:=\{1,\ldots,\Delta\}$). Furthermore, it is unknown if this Euclidean bound holds for sliding window [^1] While the result of [CJK23] works for dynamic streams, their setting is not comparable to our sliding-window streams as the deletions must be explicitly given in their dynamic streaming setting whereas the sliding-window deletes an element implicitly. . Our result nearly matches this competitive trade-off, works for sliding window, and works for general metrics which does not rely on any specific structure of Euclidean spaces [^2] Our result readily applies to Euclidean case with an additional $O(d)$ factor in space complexity, since the trivial distance oracle that returns the $\ell_{2}$ distance between two points in $\mathbb{R}^{d}$ only uses $O(d)$ space. .
@@
Our second result is a lower bound for dynamic streams, and it shows that no algorithm using small space can achieve even $\operatorname{poly}(n)$-approximation in general metric spaces. Here, the input of a dynamic stream consists of both insertions and deletions of point (IDs), and the algorithm needs to report the (estimate of the) $\operatorname{Max-Cut}$ value at the end of the stream, with respect to the inserted points that are not yet deleted. The distance oracle model is the same as in Definition 1.1, where the algorithm must provide the IDs of the point pair for querying the distance.
@@
There is no dynamic streaming algorithm that estimates the metric $\operatorname{Max-Cut}$ value within $O(\Delta/n^{1/3})$ factor using space $o(n^{1/3})$ with constant probability, where $\Delta$ is the aspect ratio and $n$ is the size of the stream.
@@
We remark that similar model for our lower bounds, which reveals the distance only upon query, were also used in e.g. [CSS16]; in fact, the model in [CSS16] is even more restricted, where the algorithm is assumed to store the IDs explicitly, and the space complexity is counted as the number of stored IDs. However, a recent work [KPSW25] introduces a stronger notion: the metric is public such that the algorithm is allowed to query any pairs, not only for the IDs in the stream (which is a subset of the full metric points). Under this model, a lower bound for diameter problem is established [KPSW25], and it is an interesting open question to establish a $\operatorname{Max-Cut}$ lower bound under the same model (which will be stronger than Theorem 1.3).
@@ 1.2.1 Upper Bounds
For insertion-only streams, our high-level idea is to (approximately) implement an importance sampling framework in general metrics, which is suggested in [CJK23] to devise a Euclidean streaming algorithm for $\operatorname{Max-Cut}$. In this framework, each data point $x\in P$ is sampled with probability roughly proportional to its total distance from all other points, denoted $q(x):=\sum_{y\in P}\mathrm{dist}(x,y)$. With a sample $S\subseteq P$ of size $\operatorname{poly}(\epsilon^{-1})$, we then evaluate $\operatorname{Max-Cut}(S)$. As in [CJK23] (restated in Lemma 3.5), this evaluation yields a $(1+\epsilon)$-approximation to $\operatorname{Max-Cut}(P)$. In our distance oracle model, evaluating $\operatorname{Max-Cut}(S)$ is straightforward since we can query the oracle to obtain all pairwise distances within $S$.
The main challenge lies in performing the importance sampling. In particular, $q(x)$ may rely on the entire stream, and it is not easy to estimate upon the arrival of $x$. To resolve this issue, our key observation is that, an approximate $\hat{q}(x)$ which estimates the sum of distances from $x$ to only the prefix until the arrival of $x$, is already a good approximation for $q(x)$, and our algorithm may be viewed as an importance sampling with respect to $\hat{q}(x)$’s.
Specifically, we propose the following modified reservoir-sampling procedure. Reservoir-sampling is a classic approach is to generate uniform samples from an insertion-only stream, and here we augment it to return importance sampling. The process maintains a single sample $s\in P$. Label the data points in $P:=\{p_{1},\ldots,p_{n}\}$ according to their order of arrival in the stream, and let $P_{i}:=\{p_{1},\ldots,p_{i}\}$ denote the prefix of the first $i$ points. When the $i$-th point $p_{i}$ is inserted we consider the prefix-based quantity $q^{\prime}(p_{i}):=\sum_{x\in P_{i}}\operatorname{dist}(p_{i},x)$. Then we replace the current sample with $p_{i}$ with probability $\frac{\hat{q}(p_{i})}{\sum_{j\leq i}\hat{q}(p_{j})}$, and do nothing otherwise.
However, it is nontrivial to maintain $\hat{q}(x)$ precisely, and we use a coreset method to approximate $\hat{q}(x)$’s. Observe that the definition of $\hat{q}(x)$ coincides with the $1$-median objective $\operatorname{med}(P_{i},x)$: for $U\subseteq V$ and $y\in V$, $\operatorname{med}(U,y):=\sum_{y^{\prime}\in U}\operatorname{dist}(y^{\prime},y)$. It is known that an $\epsilon$-coreset (see Definition 3.3) for $1$-median can be maintained efficiently in insertion-only streams (see Lemma 3.4). This result serves as a key building block for our final algorithm, and the $\epsilon$-approximation error is carefully accounted for in the analysis of Lemma 3.5.
@@
To leverage the framework of [BO10], we need to show that our insertion-only algorithm provides an approximation of a smooth function. Roughly speaking, a function $f$ is smooth, if for any $0<\epsilon<1$, there exist $0<\beta<\alpha<1$, such that if some $B\subseteq A\subseteq V$ satisfies $(1-\beta)f(A)\leq f(B)$ then $(1-\alpha)f(A\cup C)\leq f(B\cup C)$ for every $C\subseteq V$. A natural attempt is to prove the smoothness for $f=\operatorname{Max-Cut}$. However, somewhat surprisingly, $f(\cdot)=\operatorname{Max-Cut}(\cdot)$ does not satisfy this smoothness definition. To see this, consider the 1D case, and $f$ does not satisfy the condition for $A=\{1,\ldots,1\}$, then $B=\{1\}$, $C=\{0\}$, because $f(A)=f(B)=0$ but $f(A\cup C)=n>f(B\cup C)/(1-\alpha)=1/(1-\alpha)$ for any $\alpha$, where $n$ is a sufficiently large number.
We identify that the issue is caused by 0-valued $f$ (e.g., $f(A)=f(B)=0$ as in the above example). Hence, our new idea is to add a tiny term to $\operatorname{Max-Cut}$, and define $f(S):=\operatorname{Max-Cut}(S)+\frac{\epsilon}{n}|S|$ to bypass this issue, assuming the minimum non-zero distance is $1$ in the metric (otherwise one can rescale). We show that this new $f$ is indeed smooth. On the other hand, this tiny term does not add much error to $\operatorname{Max-Cut}$, so estimating $f$ suffices for estimating $\operatorname{Max-Cut}$. Indeed, if $\operatorname{Max-Cut}(P)>0$ (where $P$ is the dataset), then the added term is at most $\epsilon\operatorname{Max-Cut}(P)$, due to an assumption that the minimum non-zero distance is $1$. Hence, in this case, we can readily use the smooth histogram framework and reduce to estimating $f$ in an insertion-only stream. Finally, the other case of $\operatorname{Max-Cut}(P)=0$ can be easily detected in a sliding window and hence can be handled separately.
1.2.2 Lower Bounds
@@
At a high level, the metric of the hard input consists of many clusters such that the distance between the clusters is $\Delta\geq\Omega(n^{1/3})$ (where $n$ is the number of points), and inside the clusters the pairwise distance is $1$ (very small compared with $\Delta$), except for one special cluster of size $\Theta(n^{1/3})$, denoting whose point set as $S$. This special cluster $S$ has an outlier point whose distance to every other |S|1|S|-1 points in $S$ is some parameter $k\in\{1,\Delta\}$, thereby defining the two hard instances. The stream first sequentially inserts all points, then delete every point except for the special cluster $S$, so this special cluster $S$ is the final input dataset. If $k=1$ then $\operatorname{Max-Cut}(S)\leq O(|S|^{2})<\Delta|S|$, whereas if $k=\Delta$ then $\operatorname{Max-Cut}(S)\geq\Delta|S|$. Furthermore, any algorithm with approximation ratio $o({\Delta}/{|S|})$ must be able to tell if $k=1$ or $k=\Delta$, and this is the main claim of our proof.
Without diving into the proof details, we discuss the intuition why our input is hard in a conceptual sense. The first angle is how the deletions make it hard. Indeed, the $\operatorname{Max-Cut}$ for the entire input is easy to approximate as ignoring the special cluster, or even ignoring any constant fraction of clusters, does not change the objective by much. However, the deletion forces the algorithm to learn the structure of a tiny special cluster $S$ accurately, which is a hard task. Another angle is how we utilize the general metric and distance oracle setting, especially how this is different from the Euclidean case. In fact, the metric in our hard input can be embedded into Euclidean spaces without large distortion, hence it is not the metric itself that is hard. What makes it harder than Euclidean case, is that the algorithm can only learn the distances through a distance oracle, without the access to the vector representation in Euclidean spaces. In particular, this forces the algorithm to explicitly query the distance between the outlier point and some other point in $S$ to differentiate whether $k=1$ or $k=\Delta$ (which is crucial for approximating the $\operatorname{Max-Cut}$), and explicitly discover the outlier point is difficult. On the other hand, in Euclidean spaces one can use structures such as (space efficient) tree embedding to approximately learn the cluster structure and differentiate the outlier, without explicitly knowing the outlier point (see e.g., [CJK23]).
Due to space constraint, we defer the proof of the lower bound to Section 5.